# Eco-evolutionary constraints for the endemicity of rapidly evolving viruses

### v1

@@ Results
To shed more light into the impact of the coexistence of both evolutionary pathways on the epidemic trajectories, let us now analyze how each virus trait evolves over time. Fig. 3B shows the time evolution of the associated basic reproduction number ${\cal R}_{0}$ for each pathogen here considered. In all cases, we can observe how transmissibility evolves quicker at early than at late stages of the outbreak. This phenomenon becomes more pronounced when considering viruses with low initial infectiousness and high $D_{\lambda}$ values (see SI Appendix, Fig. S1). Consequently, our results reveal how the underlying eco-evolutionary dynamics yields an accelerated evolution in transmissibility for innocuous pathogens ($\mathcal{R}_{0}^{wt}=1.25$) compared to that experienced by more infectious pathogens, e.g. those ones with ${\cal R}_{0}^{wt}=3$.
@@
The evolution of antigenic traits can be characterized by studying the distribution of variants in the population across the antigenic space. In contrast with the evolution in transmissibility, we observe in Fig 3C how evolution in the antigenic space mainly occurs in the transition to the endemic regime after the epidemic bottleneck, reflected by the sudden enlargement in the variants distribution observed for the three pathogens. Comparing the different curves, we notice the lower the infectiousness, the higher the peak of antigenic diversity reached by the pathogen. The former result is intuitive, as those pathogens with less infectiousness should escape the immunity existing in the population more efficiently to survive in the long-term. Note that coupling both evolutionary pathways also has strong implications for the long-term evolution of the antigenic diversity. Specifically, antigenic diversity is not stabilized after the initial increase, as in Fig. 2B, but is partially lost as the virus gets more infectious to eventually reach a stable but lower value.
To shed more light into the emergent relationship between both evolutionary pathways, we compute in the Materials and Methods section an approximation for the case reproduction number ${\cal R}^{app}_{case}(t)$ representing the expected number of contagions made by one individual infected at a given day $t$ [37]. The time evolution of this quantity (Fig. 3D) confirms how antigenic and non-antigenic traits compensate one another to produce the same steady growth in the epidemic curves at late stages of the outbreak. This confirms that, despite the absence of any biological trade-off, the combination of the immune pressure in the antigenic space and the evolution in transmissibility yields universal trajectories for the evolution of pathogens in our model.
The analysis of the epidemic trajectories has revealed interesting phenomena arising from integrating evolution of antigenic and non-antigenic traits in a single formalism. In what follows, we focus on our primary research question and study the eco-evolutionary constraints for the endemicity of evolving viruses. Interestingly, Fig. 3E reveals that evolution in transmissibility drastically changes the picture observed in its absence (Fig. 2C). As stated before, evolution in transmissibility promotes virus endemicity, as the fraction of surviving realizations is always higher than in absence of such evolutionary pathway. More strikingly, in situations with strong immune pressure, i.e. low $D_{x}$ values, evolution in transmissibility gives rise to a non-mononotic behavior of the virus endemicity with the infectiousness of the wild-type variant. For instance, the curve corresponding to $D_{x}=0.015$ shows that viruses with an initial reproduction number of $\mathcal{R}_{0}^{wt}\simeq 1$ are more prone to persist in the population than others with $\mathcal{R}_{0}^{wt}\simeq 2$. Note that this unexpected behavior dissappears when immune pressure loses relevance, i.e. for high $D_{x}$ values or high $D_{\lambda}$ values (see SI Appendix, Fig. S1D), retrieving the monotonic behavior typically reported in the literature.
For the sake of completeness, we also consider a scenario where the virus evolution is bounded by a transmission-recovery trade-off accelerating the recovery rate of hosts for more transmissible pathogens (see SI Appendix, Fig. S2). Biologically, the latter assumption is justified as an enhanced immune response for pathogens replicating more efficiently within the host [38]. The results here reported remain robust as long as the relative increase in the recovery rate is negligible. Nonetheless, when the trade-off gains relevance, the non-monotonic behavior is lost. In this case, the initial evolution in transmissibility shortens the characteristic time scale of outbreaks, thus hampering immune escape. Moreover, under the transmission-recovery trade-off, viruses are able to find evolutionary stable strategies (ESS), thus eventually slowing down the evolution in transmissibility and therefore the chances for pathogens to become endemic. More future research will be relevant to determine how the phenomena here reported is sensitive to other biological trade-offs.
.4 Epidemiological and evolutionary time scales shape endemicity
Our results from stochastic simulations reveal that the interplay between immune pressure and evolution in transmissibility favors the endemicity of weakly infectious pathogens. To understand the origin of such phenomenon, we first realize that, in absence of transmissibility evolution (Fig. 2C), there is a critical value of the basic reproduction number ${\cal R}^{wt,C}_{0}(D_{x})$ below which extinction of the virus always occurs. This critical value is lower for viruses with efficient immune escape, i.e. high $D_{x}$ values. Inspecting visually the critical values in the endemicity curves for viruses now evolving their transmissibility (Fig. 3E), we notice that the unexpected non-monotonic behavior involves pathogens whose infectiousness falls below ${\cal R}^{wt,C}_{0}(D_{x})$. Conversely, the region ${\cal R}_{0}^{wt}>{\cal R}^{wt,C}_{0}(D_{x})$ retrieves the monotonically increasing behavior with the infectiousness of the wild-type variant. From this qualitative analysis, it becomes clear that the non-monotonic behavior previously described arises from the evolutionary dynamics undergone for viruses with ${\cal R}_{0}^{wt}<{\cal R}^{wt,C}_{0}(D_{x})$ to reach that critical value before becoming extinct.
To further support our intuition, we seek to analytically derive the eco-evolutionary constraints determining the endemicity of rapidly evolving viruses. Unfortunately, the underlying stochasticity of both evolutionary and epidemiological processes makes the analytical treatment of our model quite cumbersome, yet some theoretical results can be obtained in specific circumstances [32, 19]. To overcome this issue, we move to a more simplified eco-evolutionary framework, where we assume that both transmissibility and antigenic position evolve deterministically at a constant pace. Hence, for each time step $t$, there is a single variant circulating in the population, characterized by $x(t)=\widetilde{D}_{x}t$ and $\lambda(t)=\widetilde{D}_{\lambda}t$, where $\widetilde{D}_{m}$ represent the deterministic speed of evolution of trait $m$. Note that this framework cannot reproduce some of our previous findings, such as the accelerated evolution for the most innocuous viruses at initial stages of epidemic outbreaks. Notwithstanding its relevance, we will prove that such acceleration plays a minor role in the emergence of the non-monotonic behavior found for viral endemicity.
Introducing deterministic evolutionary pathways considerably eases the mathematical characterization of the underlying eco-evolutionary dynamics. In absence of evolution in transmissibility, we can assume that endemicity is determined by the prevalence of the disease at equilibrium, denoted in what follows by $\rho_{I}^{*}$. Using a travelling wave approach and considering a scenario with strong immune pressure, we show in the Materials and Methods section that the epidemic prevalence at equilibrium reads:
$\rho_{I}^{*}\simeq\frac{2\mathcal{R}^{wt}_{0}D_{x}}{\mu\pi}\ .$ (2)
If the prevalence $\rho_{I}^{*}$ is small enough, stochastic fluctuations associated with contagions and recoveries drive the system to the absorbing state and the virus becomes extinct. Considering the existence of a critical value $\rho_{I}^{*,C}$ to avoid the former scenario, we obtain that the minimum speed of evolution in the antigenic space that a virus requires to become endemic, denoted by $\widetilde{D}_{x}^{C}$, fulfills:
$\widetilde{D}_{x}^{C}=\frac{\mu a(N)}{\mathcal{R}^{wt}_{0}}\ ,$ (3)
where $a(N)=\pi\rho_{I}^{*,C}/2$ captures the aforementioned stochastic effects and depends on the population size $N$. In what follows, we will consider $a$ as a scaling factor to be fitted. To show the validity of this approach, we represent in Fig. 4A the fraction of realizations becoming endemic as a function of both $\widetilde{D}_{x}$ and $\mathcal{R}^{wt}_{0}$, proving that Eq. 3 captures the eco-evolutionary constraints allowing the persistence of viruses. As in the stochastic case, in absence of evolution in transmissibility, endemicity monotonically increases with $\mathcal{R}^{wt}_{0}$. Eq. 3 also captures the critical surfaces for pathogens with different $\mu$ values (SI Appendix, Fig. S3), showing that viruses with shorter infectious windows should evolve more quickly in their antigenic space to achieve endemicity. Likewise, we find that both $\widetilde{D}_{x}^{C}$ and, henceforth, $a$ are decreasing functions of the population size $N$, which is an expected result as stochastic fluctuations are less relevant in large populations (SI Appendix, Fig. S4).
Figure 4: Endemicity of viruses under deterministic evolution. A)-D): Fraction of endemic realizations (color code) as a function of the basic reproduction number of the wild-type variant $\mathcal{R}^{wt}_{0}$ and the speed of evolution in the antigenic space $\widetilde{D}_{x}$. The white solid line shows the theoretical estimation of the critical immunity escape value $\widetilde{D}^{C}_{x}$ delimiting the region $f_{endemic}=1$. Such quantity is obtained by setting $a=1/12$ and $b=5$ in Eq. 4. The values considered for the speed of evolution in transmissibility are: (A) $\widetilde{D}_{\lambda}=0$, (B) $\widetilde{D}_{\lambda}=4\cdot 10^{-5}$, (C) $\widetilde{D}_{\lambda}=8\cdot 10^{-5}$ and (D) $\widetilde{D}_{\lambda}=1.2\cdot 10^{-4}$. In all panels, the fraction of endemic realizations is obtained by performing 500 epidemic outbreaks and computing those persisting in the population after $t=1000$ days. The rest of model parameters are the same as in Fig. 2.
Now we modify Eq. 3 to capture how evolution in transmissibility shapes the endemicity of viruses. As the transmissibility evolves in an unbounded way, stochastic fluctuations during the epidemic bottleneck are the most relevant mechanism for the extinction of the virus. At that point, the virus infectiousness is determined by its gain in transmissibility during the first epidemic wave. Unfortunately, given the non-linearity of the SIR equations, the duration of such phase, denoted by $\tau_{wave}$, cannot be obtained analytically. Nonetheless, heuristically, we can argue that the characteristic time scale of an epidemic wave is proportional to the time to reach the peak of infected individuals, $t_{peak}$, which has been recently estimated very accurately [39]. Introducing the latter assumption into Eq. 3 yields:
where the expression for $t_{peak}$ is provided in the Materials and Methods section and $b$ is another scaling factor to be fitted to obtain the characteristic time scale of the disease, i.e. $\tau_{wave}=bt_{peak}$.
Figs. 4B-D show how the latter equation fairly captures the non-monotonic behavior of the endemicity with ${\mathcal{R}^{wt}_{0}}$ for different values of the speed of evolution in transmissibility $D_{\lambda}$. Note that there are some deviations for $\mathcal{R}_{0}\simeq 1$ and high $D_{\lambda}$ values, as the expression for $t_{peak}$ neglects the underlying evolution of the virus until reaching the epidemic peak, which is especially relevant in that region of the parameters space.
More importantly, Eq. 4 explains why viruses with intermediate infectiousness are less likely to become endemic. Namely, these viruses are not infectious enough to persist in the population without evolving their transmissibility but their generated first epidemic wave is not long enough to reach the critical transmissibility value ${\mathcal{R}_{0}^{wt,C}(D_{x})}$. Conversely, the nonlinear inverse dependence of $t_{peak}$ on $\mathcal{R}^{wt}_{0}$ (Eq. 11) makes the duration of the first epidemic wave much longer for pathogens with lower initial infectiousness, then giving them the chance to evolve before becoming extinct.
Discussion
Understanding the evolutionary forces governing how infectious diseases transition to become endemic in real populations represents a timely research question [40]. Here we propose a minimal eco-evolutionary framework to characterize how the evolution in both antigenic and non-antigenic traits shape the course and fate of epidemic outbreaks. Our results from stochastic simulations have first shown that evolution is not constant in both dimensions across time. Instead, the interplay between ecological and epidemiological processes accelerates the evolution in transmissibility at early stages of the outbreak while prioritizing the antigenic escape once the first epidemic wave is over. The accelerated evolution in fitness at early generations of RNA viruses has already been reported empirically in culture cells [41] and characterized theoretically through Fokker-Planck equations [42]. Likewise, the shift in the role of different evolutionary pathways has also been observed in the SARS-CoV-2 virus. Analysis of genomic data [43] reveals that intrinsic selection has become less relevant, with antigenic fitness gaining prominence throughout the COVID-19 pandemic.
Integrating the evolution of both antigenic and non-antigenic traits drastically changes the eco-evolutionary constraints shaping the endemicity of rapidly evolving viruses. Namely, we have observed how immune escape alone predicts that more infectious pathogens are more likely to become endemic. Conversely, in presence of evolution in transmissibility, viruses with low basic reproduction number, i.e ${\mathcal{R}^{wt}_{0}}\simeq 1$, are more likely to persist in the population that those with intermediate infectiousness, e.g. ${\mathcal{R}^{wt}_{0}}\simeq 2$. To characterize analytically this behavior, we have developed a simplistic model assuming that virus traits evolve deterministically. Our analysis reveals that the non-monotonic behavior emerges because the latter viruses generate much shorter epidemic outbreaks, thus not being able to evolve their transmissibility to reach the critical value needed to become endemic. Interestingly, the non-monotonic behavior of the endemicity of epidemic outbreaks has been already reported for other variants of the SIR model including demographic births and deaths [44, 45].
From a practical point of view, our findings represent a cautionary tale for the implementation of control strategies to manage epidemic scenarios. While the short-term beneficial effects of non-pharmaceutical interventions [46, 47] cannot be contested, determining their long-term impact on the modified epidemic trajectories still represents a more complex and intricate problem with growing attention in the scientific community [48, 49, 50]. Along this line, our results show that controlling an epidemic scenario by reducing the local effective reproduction number ${\cal R}$ might render unforeseen long-term consequences for disease control. In particular, our results show that mitigating an outbreak by flattening the epidemic curves induces a stronger selection pressure on the virus and increases the risk for their persistence in the population, due to the extended duration of the epidemic outbreak. Likewise, other works have reported that vaccination of the population also alters the antigenic space harboring the different variants of the virus [51, 52], thus potentially shaping the evolutionary strategies observed at the population level [53]. Moving from local to global interventions, the evolutionary fingerprint of travel bans and the subsequent geographical genetic isolation of different areas [54] remains to be solved; phylogeographic analyses [55] combined with multiscale epidemic frameworks [56] will be needed to address this question in future works.

### v2

@@ Results
Let us now analyze how each virus trait evolves over time. Fig. 3B shows the time evolution of the associated basic reproduction number ${\cal R}_{0}$ for each pathogen here considered. For each realization, we consider the mean infectiousness of the circulating variants at a given time step $t$, $\bar{\lambda}(t)$, to compute this indicator, yielding ${\cal R}_{0}(t)=\bar{\lambda}(t)k/\mu$. In all cases, we can observe how infectiousness evolves quicker at early than at late stages of the outbreak. This phenomenon becomes more pronounced when considering viruses with low initial infectiousness and high $D_{\lambda}$ values (see Appendix B). Consequently, our results reveal how the underlying eco-evolutionary dynamics yields an accelerated evolution in infectiousness for innocuous pathogens ($\mathcal{R}_{0}^{wt}=1.25$) compared to that experienced by more infectious pathogens, e.g. those ones with ${\cal R}_{0}^{wt}=3$.
Figure 3: Eco-evolutionary dynamics under evolution of antigenic and non-antigenic traits. A)-D): Time evolution of different epidemiological quantities and virus traits in endemic epidemic outbreaks. The quantities shown correspond to: (A) fraction of infected population $\rho_{I}$, (B) basic reproduction number ${\cal R}_{0}$, (C) variance of the distribution of strains across the antigenic space $\sigma^{2}_{x}$ and (D) an estimation of the case reproduction number $R^{app}_{case}$ (see Appendix C). In all these panels, line color corresponds to the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$.The symbol $\langle\cdot\rangle$ denotes that each curve is the result of averaging the individual curves of all endemic realizations observed after simulating 10001000 epidemic outbreaks for each ${\cal R}_{0}$ value. The speeds of evolution in the infectiousness and antigenic spaces are set to $D_{\lambda}=0.0003$ and $D_{x}=0.015$ respectively. E): Endemicity $f_{endemic}$ of the virus as a function of the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$. Line color here represents the speed of evolution in the antigenic space $D_{x}$. Solid (dotted) lines represent the values found in presence (absence) of evolution in the infectiousness space by setting $D_{\lambda}=0.0003$ ($D_{\lambda}=0$). In all the panels, the rest of epidemiological parameters are the same as in Fig. 2.
The evolution of antigenic traits can be characterized by the distribution of strains across the antigenic space. In contrast with the evolution in infectiousness, we observe in Fig 3C how evolution in the antigenic space mainly occurs in the transition to the endemic regime after the epidemic bottleneck, reflected by the sudden enlargement in the variants distribution observed for the three pathogens. Comparing the different curves, we notice the lower the infectiousness, the higher the peak of antigenic diversity reached by the pathogen. This result is intuitive, as those pathogens with less infectiousness should accumulate more antigenic diversity to survive in the population. Note that coupling both evolutionary pathways also has strong implications for the long-term evolution of the antigenic diversity. Specifically, antigenic diversity is not stabilized after the initial increase, as in Fig. 2B, but is partially lost as the virus gets more infectious to eventually reach a stable but lower value.
To characterize the emergent relationship between both evolutionary pathways, in Appendix C we compute a theoretical approximation for the case reproduction number ${\cal R}^{app}_{case}(t)$ representing the expected number of contagions made by one individual infected at a given day $t$ [37]. The time evolution of this quantity (Fig. 3D) illustrates how antigenic and non-antigenic traits compensate one another to produce the same steady growth in the epidemic curves at late stages of the outbreak. This confirms that, despite the absence of any biological trade-off, the combination of immune pressure in the antigenic space and evolution in infectiousness yields universal trajectories for the evolution of pathogens in our model.
We now focus on our primary research question and study the eco-evolutionary constraints for the endemicity of evolving viruses. Interestingly, Fig. 3E reveals that evolution in infectiousness drastically changes the picture observed in its absence (Fig. 2C). As stated before, evolution in infectiousness promotes viral endemicity, as the fraction of surviving realizations is always higher than in absence of such evolutionary pathway. More strikingly, in situations with little antigenic diversity, i.e. low $D_{x}$ values, evolution in infectiousness gives rise to a non-monotonic behavior of the virus endemicity with the infectiousness of the wild-type variant. For instance, the curve corresponding to $D_{x}=0.015$ shows that viruses with an initial reproduction number of $\mathcal{R}_{0}^{wt}\simeq 1$ are more prone to persist in the population than others with $\mathcal{R}_{0}^{wt}\simeq 2$. Note that this unexpected behavior disappears when immune pressure loses relevance, i.e. for high $D_{x}$ values or high $D_{\lambda}$ values (see Appendix B), retrieving the monotonic behavior typically reported in the literature.
For the sake of completeness, we also consider a scenario where viral evolution is bounded by a transmission-recovery trade-off accelerating the recovery rate of hosts for more transmissible pathogens (see Appendix D for more details). Biologically, the latter assumption is justified as an enhanced immune response for pathogens which replicate more efficiently within the host [38]. Our results show that the interplay between evolution and endemicity remains robust as long as the relative increase in the recovery rate is negligible. Nonetheless, when the trade-off gains relevance, the non-monotonic behavior is lost. In this case, the initial evolution in infectiousness shortens the characteristic time scale of outbreaks, thus limiting immune escape. Moreover, under the transmission-recovery trade-off, viruses are able to find evolutionarily stable strategies (ESS), thus eventually slowing down the evolution in infectiousness and therefore the chances for pathogens to become endemic. More future research will be relevant to determine how the phenomena here reported are sensitive to other biological trade-offs.
Evolution of infectiousness in the first epidemic wave explains the non-monotonic endemicity with ${\cal R}_{0}^{wt}$
Our results from stochastic simulations reveal that the interplay between immune pressure and evolution in infectiousness favors the endemicity of weakly infectious pathogens. To understand the origin of such phenomenon, we first realize that, in absence of infectiousness evolution (Fig. 2C), there is a critical value of the basic reproduction number ${\cal R}^{wt,C}_{0}(D_{x})$ below which viruses become extinct. This critical value is lower for viruses with efficient immune escape, i.e. high $D_{x}$ values. Inspecting visually the critical values in the endemicity curves for viruses now evolving their infectiousness (Fig. 3E), we notice that the unexpected non-monotonic behavior involves pathogens whose basic reproduction number falls below ${\cal R}^{wt,C}_{0}(D_{x})$. Conversely, the region ${\cal R}_{0}^{wt}>{\cal R}^{wt,C}_{0}(D_{x})$ retrieves the monotonically increasing behavior with the infectiousness of the wild-type variant. From this qualitative analysis, it becomes clear that the non-monotonic behavior previously described arises from the evolutionary dynamics undergone by viruses with ${\cal R}_{0}^{wt}<{\cal R}^{wt,C}_{0}(D_{x})$ to reach that critical value before becoming extinct.
To further support our intuition, we seek to analytically derive the eco-evolutionary constraints determining the endemicity of rapidly evolving viruses. Let us first neglect the evolution in infectiousness. In this case, the state of our system at time $t$ is characterized by the distributions, $\rho_{I}(x,t)$ ($\rho_{R}(x,t)$), defined as the fraction of the population in the infected (recovered) state carrying (with immunity to) the strain $x$ at time $t$. Considering antigenic escape and reinfection events, the time evolution of the latter is given by [19, 32]:
$\frac{d\rho_{R}(x,t)}{dt}$ $=\mu\rho_{I}(x,t)$
$-\lambda k\rho_{R}(x,t)\int_{x}^{\infty}\rho_{I}(x^{\prime},t)\left(1-e^{-(x^{\prime}-x)}\right)dx^{\prime}\;,$ (2)
where the first term corresponds to the recovery of infected individuals and the second term encodes the loss of recovered individuals due to reinfection events by variants with different antigenic features.
Unfortunately, the underlying stochasticity of both evolutionary and epidemiological processes makes the analytical treatment of this model quite cumbersome, yet some theoretical results can be obtained in specific circumstances [32, 19]. To overcome this issue, we move to a more simplified eco-evolutionary framework, where we assume that both infectiousness and antigenic position grows linearly with time following their evolution inside infected hosts. Note that, in doing so, we cannot capture some of our previous findings, such as the accelerated evolution of the most innocuous viruses at early stages of the epidemic outbreak. Nonetheless, this constant evolution, yet unrealistic, captures the increasing infectiousness and antigenic position due to the selection driven by epidemiological processes.
Figure 4: Endemicity of viruses under deterministic evolution. A)-D): Fraction of endemic realizations $f_{endemic}$ as a function of the basic reproduction number of the wild-type variant $\mathcal{R}^{wt}_{0}$ and the speed of evolution in the antigenic space $\widetilde{D}_{x}$. The white solid line shows the theoretical estimation of the critical immunity escape value $\widetilde{D}^{C}_{x}$ delimiting the region $f_{endemic}=1$. Such quantity is obtained by setting $a=1/12$ and $b=5$ in Eq. 6. The values considered for the speed of evolution in infectiousness are: (A) $\widetilde{D}_{\lambda}=0$, (B) $\widetilde{D}_{\lambda}=4\cdot 10^{-5}$, (C) $\widetilde{D}_{\lambda}=8\cdot 10^{-5}$ and (D) $\widetilde{D}_{\lambda}=1.2\cdot 10^{-4}$. In all panels, the fraction of endemic realizations is obtained by performing 500 epidemic outbreaks and computing those persisting in the population after $t=1000$ days. The rest of model parameters are the same as in Fig. 2.
The constant evolution of viral traits eases the mathematical formulation of the model, as all the infected individuals at a given time step $t$ share the same strain $c$, characterized by the antigenic position $x_{c}(t)=\widetilde{D}_{x}t$, where $\widetilde{D}_{x}$ refers to the (constant) speed of antigenic evolution. Consequently, the spatiotemporal distribution of infected population across the antigenic space reads $\rho_{I}(x,t)=\rho_{I}(t)\delta\left(x-x_{c}(t)\right)$. Using this expression and considering a travelling wave approach, we can rewrite Eq. (2) as:
where $\tau=(x_{c}-x)/\widetilde{D}_{x}$ refers to the time elapsed since agents’ recovery. To estimate the eco-evolutionary constraints shaping the endemicity of virus, we focus on the prevalence of the disease at the epidemic bottleneck after the first epidemic wave, denoted by $\rho_{I}^{bottleneck}$. When the latter prevalence is small enough, stochastic fluctuations in recovery events drive the system to the absorbing state and the virus becomes extinct, despite its potential to generate another outbreak. Assuming that $\rho_{I}^{bottleneck}$ remains constant in the bottleneck in a critical scenario, we can derive its value from Eq. (3) as detailed in Appendix E, obtaining:
$\rho_{I}^{bottleneck}\simeq\frac{2\mathcal{R}^{wt}_{0}\widetilde{D}_{x}}{\mu\pi}\ .$ (4)
Considering a critical $\rho_{I}^{bottleneck,C}$ value preventing the stochastic fadeout of simulations, we obtain that the minimum speed of evolution in the antigenic space that a virus requires to become endemic, denoted by $\widetilde{D}_{x}^{C}$, fulfills:
$\widetilde{D}_{x}^{C}=\frac{\mu a(N)}{\mathcal{R}^{wt}_{0}}\ ,$ (5)
where we assume $a(N)=\pi\rho_{I}^{bottleneck,C}/2$ to depend on the population size $N$, as smaller populations are, in principle, more vulnerable to finite size effects.
In what follows, we will consider $a$ as a scaling factor to be fitted. To show the validity of this approach, we represent in Fig. 4A the fraction of realizations becoming endemic as a function of both $\widetilde{D}_{x}$ and $\mathcal{R}^{wt}_{0}$, proving that Eq. (5) captures the eco-evolutionary constraints allowing the persistence of viruses. As in the original model, in absence of evolution in infectiousness, endemicity monotonically increases with $\mathcal{R}^{wt}_{0}$. Eq. (5) also captures the critical surfaces for pathogens with different $\mu$ values (Appendix F), showing that viruses with shorter infectious windows should evolve more quickly in their antigenic space to achieve endemicity. Likewise, we find that both $\widetilde{D}_{x}^{C}$ and, henceforth, $a$ are decreasing functions of the population size $N$, which is an expected result as stochastic fluctuations are less relevant in large populations (Appendix F).
Now we modify Eq. (5) to capture how evolution in infectiousness shapes the endemicity of viruses. Under constant evolution, the infectiousness of the circulating strain $c$ at time $t$ reads $\lambda_{c}(t)=\lambda^{wt}+\widetilde{D}_{\lambda}t$. At the epidemic bottleneck, the infectiousness of the virus is determined by the increments accumulated from evolution during the first epidemic wave. Unfortunately, given the non-linearity of the SIR equations, the duration of this phase, denoted by $\tau_{wave}$, cannot be obtained analytically. However, heuristically, we can argue that the characteristic time scale of an epidemic wave is proportional to the time to reach the peak of infected individuals, $\tau_{peak}$, which has been recently estimated [39]. Taking into account the evolution in the reproduction number during the first epidemic wave, Eq. 5 turns into:
where the expression for $t_{peak}$ is provided in Appendix G and $b$ is another scaling factor to be fitted to obtain the characteristic time scale of the disease, i.e. $\tau_{wave}=b\tau_{peak}$.
Figs. 4B-D show how the latter equation fairly captures the non-monotonic behavior of the endemicity with $\mathcal{R}^{wt}_{0}$} for different values of the speed of evolution in infectiousness $D_{\lambda}$. Note that there are some deviations for $\mathcal{R}_{0}\simeq 1$, as the expression for $t_{peak}$ neglects the underlying evolution of the virus until reaching the epidemic peak, which is especially relevant in that region of the parameters space.
More importantly, Eq. 6 explains why viruses with intermediate infectiousness are less likely to become endemic. For these viruses, ${\cal R}_{0}^{wt}<{\mathcal{R}_{0}^{wt,C}(D_{x})}$ but also ${\cal R}_{0}^{wt}+bD_{\lambda}t_{peak}(\mathcal{R}^{wt}_{0},\mu)k/\mu<{\mathcal{R}_{0}^{wt,C}(D_{x})}$, given the short duration of their first epidemic waves. Conversely, the nonlinear inverse dependence of $\tau_{peak}$ on $\mathcal{R}^{wt}_{0}$ (Appendix G) makes the duration of the first epidemic wave much longer for pathogens with lower initial infectiousness, thus allowing them to evolve to reach the critical infectiousness before becoming extinct. In Appendix F, we analyze in depth the epidemic trajectories obtained from the simulations, finding that viruses with intermediate infectiousness present stronger epidemic bottlenecks and are not able to accumulate neither enough infectiousness nor enough antigenic diversity to generate secondary outbreaks consistently.
Figure 5: Trade-off between short-term and long-term benefits of control policies. A): Time evolution of the epidemic prevalence $\rho_{I}$. Line color denotes the duration of the control policies $\tau_{control}$. B): Distribution of epidemic peak as a function of $\tau_{control}$. C): Time evolution of the antigenic diversity $\sigma_{x}^{2}$. Line color denotes the value of $\tau_{control}$. D): Distribution of antigenic diversity at the epidemic bottleneck across endemic realizations, $(\sigma^{2}_{x})^{bottleneck}$, as a function of $\tau_{control}$. We define the epidemic bottleneck as the point with minimum incidence following the first epidemic wave. In panels A) and C) thick lines represent average values across endemic realizations whereas thin lines correspond to single realizations. In panels B) and D), the dot shows the mean of the distribution whereas the whiskers denote its IQR. In all these panels, the (constant) speed of antigenic evolution is set to $\widetilde{D}_{x}=0.0003$. E): Fraction of endemic realizations as a function of $\tau_{control}$. Line color encodes the value of $\widetilde{D}_{x}$. The criterion used to classify a realization as endemic is the same as the one used in Fig. 4. In all panels, we do not consider evolution in infectiousness, i.e. $\widetilde{D}_{\lambda}=0$ and the epidemiological parameters are the same as in Fig. 2.
Control policies in the first epidemic wave affects the long-term behavior of epidemics
Our results reveal that both the evolved infectiousness and the antigenic diversity generated during the first epidemic wave are crucial to understand viral endemicity. To single out the relevance of antigenic diversity, we now consider viruses not evolving their infectiousness, i.e. $\widetilde{D}_{\lambda}=0$ and tune the duration of the first epidemic wave by introducing control measures flattening epidemic curves. These control policies are activated when a fraction $\theta=0.01$ of the population is infected and lifted after $\tau_{control}$ time steps. Among the different choices, we consider that control policies halve the number of contacts of the population, passing from $k=10$ to $k=5$ contacts. Consequently, the reproduction number of the disease is halved with respect to the uncontrolled scenario during the controlled outbreak.
Fig.5A shows the infection curves for a virus with ${\cal R}_{0}^{wt}=2$ and $\widetilde{D}_{x}=0.003$. As expected, halving the reproduction number renders short-term benefits, for the epidemic peak is reduced as the duration of the policies is increased (see Fig.5B). Nonetheless, the effects of such policies on the antigenic diversity $\sigma_{x}^{2}$ are more complex (Fig.5C). In particular, Fig.5D reveals that increasing the duration of control policies gives rise to more antigenic diversity at the epidemic bottleneck. This greater antigenic diversity enables more reinfection events and increases the fraction of endemic realizations (Fig.5E), revealing that flattening epidemic curves might enhance the long-term persistence of viruses.
Discussion
Understanding the evolutionary forces governing how infectious diseases transition to become endemic in real populations represents a timely research question [40]. Here we propose a minimal eco-evolutionary framework to characterize how the evolution in both antigenic and non-antigenic traits shape the course and fate of epidemic outbreaks. Our results from stochastic simulations have first shown that evolution is not constant in both dimensions across time. Instead, the interplay between ecological and epidemiological processes accelerates the evolution in infectiousness at early stages of the outbreak while favoring antigenic diversification once the first epidemic wave is over. This phenomenon occurs because the existence of immune pressure in the population might render mutations reducing infectiousness but changing antigenic properties of viruses beneficial, thus slowing down the overall increase of viral infectiousness. Note that this result cannot be found in Susceptible-Infected-Recovered-Susceptible models with constant waning immunity rates. Empirically, the accelerated evolution in fitness at early generations of RNA viruses has already been reported in culture cells [41] and characterized theoretically through Fokker-Planck equations [42, 43]. Likewise, genomic surveillance data of the SARS-CoV-2 [44] reveals the shift from intrinsic selection to antigenic diversification during the COVID-19 pandemic. Yet both evolutionary pathways are in principle uncorrelated, our simulations show that immune pressure in the population also acts a selection force intertwining them, giving rise to universal long-term evolutionary trajectories.
Integrating the evolution of both antigenic and non-antigenic traits drastically changes the eco-evolutionary constraints shaping the endemicity of rapidly evolving viruses. Namely, we have observed how immune escape alone predicts that more infectious pathogens are more likely to become endemic. Conversely, in presence of evolution in infectiousness, viruses with low basic reproduction number, i.e ${\mathcal{R}^{wt}_{0}}\simeq 1$, are more likely to persist in the population than those with intermediate infectiousness, e.g. ${\mathcal{R}^{wt}_{0}}\simeq 2$. To characterize analytically this behavior, we have developed a simplistic model assuming a constant evolution in viral traits. Our analysis reveals that the non-monotonic behavior emerges because the latter viruses generate much shorter epidemic outbreaks, thus not being able to accumulate neither enough viral diversity nor enough infectiousness to persist in the population.
The non-monotonic behavior of viral endemicity with respect to the basic reproduction number has already been reported for a SIR model including birth and death dynamics [45, 46]. In this model, the replenishment of the depleted pool of susceptible individuals occurs due to the replacement of dead population by newborn susceptible ones. Interestingly, the non-monotonic behavior emerges theoretically when the duration of the first epidemic wave becomes comparable with the lifetime of individuals. As argued in [45], this non-monotonic behavior driven by birth and death dynamics cannot be observed for real diseases, as their typical time scales much shorter than the underlying population dynamics. Sharing the same physical principles, our study provides a more plausible origin of this phenomenon rooted in the replenishment of susceptible population due to antigenic diversification and evolution in infectiousness during the first epidemic wave. Further empirical work will be needed to measure the evolutionary rates of viruses and check whether this non-monotonic behavior is a mere theoretical result or instead can play an important role in real outbreaks.